OpenStudy (anonymous):

integral sec^2(6x) dx integral sec^2(6x) dx @Mathematics

OpenStudy (anonymous):

tan(6x)/6 steps to follow.

OpenStudy (anonymous):

u=6x du=6dx (1/6)sec(u)^2 integral of sec(u)^2 = tan(u) plug it in. (1/6)*tan(6x) + C

OpenStudy (agreene):

For the integrand sec^2(6 x), substitute u = 6 x and du = 6 dx: = 1/6 integral sec^2(u) du The integral of sec^2(u) is tan(u): = (tan(u))/6+constant Substitute back for u = 6 x: = 1/6 tan(6 x)+constant